geerky42 (geerky42):

Knowing that \(2\sin x + \cos x = 1\), find \(\tan x\).

OpenStudy (raden):

x=0, the first guess :)

geerky42 (geerky42):

Are there any more solutions?

OpenStudy (raden):

2sinx+cosx=1 divided by cosx to both sides giving us 2tanx + 1 = secx squares to both sides : 4tan^2 x + 4tanx + 1 = sec^2 x 4tan^2 x + 4tanx + 1 = 1 + tan^2 x 3tan^2 x + 4tanx = 0 tanx(3tanx + 4) = 0

OpenStudy (raden):

take the zeroes of all factors

OpenStudy (raden):

btw, u should give intervals for x

OpenStudy (anonymous):

\[2sinx=1-cosx\] after applying half-angle formulae, \[2*2\sin \frac{x}{2}\cos \frac{x}{2}=2\sin^{2} \frac{x}{2}\] \[2\cos \frac{x}{2}=\sin \frac{x}{2}\] \[\tan \frac{x}{2}=2\] apply this for tanx. \[\tan(\frac{x}{2}+ \frac{x}{2}) = \frac {\tan(\frac {x}{2})+\tan(\frac {x}{2})}{1-\tan(\frac {x}{2})\tan(\frac {x}{2})}\]

hartnn (hartnn):

one more alternate method, for \(a\sin x+b\cos x=c\), divide both sides by \(\sqrt{a^2+b^2}\)